Wildhorse Resources Company is currently an all-equity firm with a WACC of 17 percent and a 40 percent marginal tax rate. Wildhorse wants to move to a capital structure with $300 million of debt outstanding at an interest rate of 6 percent and a market value of equity equal to $900 million outstanding. Using M&M Proposition 2 with taxes, Equation 16.5, what is the expected return on equity at the new capital structure? (Round answer to 1 decimal place, e.g. 17.5%.) Expected return on equity enter the expected return on equity in percentages rounded to 1 decimal place %
Answer #1

rE = rA + [(D/E) * (1 - t) * (rA - rD)]

= 0.17 + [(300/900) * (1 - 0.40) * (0.17 - 0.06)]

= 0.17 + [0.3333 * 0.60 * 0.11] = 0.17 + 0.022 = 0.192, or 19.2%